OpenStudy (anonymous):

Last week, Eva spent 7 and 3/8 hours on homework. This week, she spent 10 5/8 hours. How many more hours did she spend this week?

OpenStudy (one098):

So, what you need to do is subtract. \[7\frac{ 3 }{ 8 }-10\frac{ 5 }{ 8 }\] First subtract the whole numbers. So 7-10. Then subtract 3-5 and you just keep the denominator.

OpenStudy (anonymous):

\[\frac{85}{8}-\frac{59}{8}=\frac{13}{4}=3+\frac{1}{4} \]